# Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models

### v1

@@ 3.2 Positive Adaptation
Positive Textual Adapter. As shown in Eq. (2), CLIP can make zero-shot predictions utilizing a textual cache, which stores the textual features from positive text descriptions (e.g., “A photo of a {CLASS}”). To avoid ambiguity, we denote this cache as $\mathsf{T}_{\mathtt{cache}}^{+}=[f_{t_{1}}^{+}\,f_{t_{2}}^{+}\,\cdots\,f_{t_{C}}^{+}]^{\top}\in\mathbb{R}^{C\times d}$. With a small set of annotated training images, we can improve this cache-based classifier by enhancing the inter-modal alignment between textual features and the few-shot image features from the same class. Specifically, we follow TaskRes [75] to achieve this by introducing a group of learnable parameters $\mathcal{R}_{\mathsf{T}}^{+}\in\mathbb{R}^{C\times d}$. By incorporating these parameters with the textual cache $\mathsf{T}_{\mathtt{cache}}^{+}$ via element-wise addition in a residual approach, it integrates task-specific knowledge and enhances the inter-modal classifier:
$\mathsf{T}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{T}}^{+}\right),\quad\mathcal{S}_{\mathsf{T}}^{+}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{+\quad\top}\in\mathbb{R}^{1\times C}.$ (3)
Here, $\mathsf{Normalize}$ denotes the L2-normalization applied to each row of the matrix.
Positive Visual Adapter. Inspired by Tip-Adapter [79], we further extend the recognition capability of the CLIP model by constructing a visual cache-based classifier, which operates by measuring intra-modal similarities between the input image feature and few-shot image features. Given a $C$-class $K$-shot training dataset from a specific downstream task, we store all $CK$ image features in a precomputed visual cache, denoted as $\mathsf{V}_{\mathtt{cache}}^{+}=[f_{v_{1}}^{(1)}\,f_{v_{1}}^{(2)}\,\cdots\,f_{v_{C}}^{(K)}]^{\top}\in\mathbb{R}^{CK\times d}$. Their one-hot labels are also correspondingly recorded in $L\in\mathbb{R}^{CK\times C}$. Given an image feature $f_{\mathcal{I}}$ to be classified, we calculate its image-image affinities $\mathcal{A}^{+}$ with all the few-shot training images, then multiplied by their corresponding one-hot labels $L$ to obtain the classification logit $\mathcal{S}_{\mathsf{V}}^{+}$:
$\mathcal{A}^{+}=\exp\left(-\beta\left(1-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\right)\in\mathbb{R}^{1\times CK},\quad\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}L\in\mathbb{R}^{1\times C},$ (4)
where $\alpha$ represents a balance factor and $\beta$ denotes a modulating hyper-parameter. Empirically, the value of $\alpha$ should be increased in cases of larger domain shifts, where more domain-specific few-shot knowledge is needed for adaptation.
Symmetrically, we also introduce a set of learnable parameters $\mathcal{R}_{\mathsf{V}}^{+}\in\mathbb{R}^{C\times d}$, which are broadcast to $\mathbb{R}^{CK\times d}$ and added to the positive visual cache:
@@
$\mathsf{V}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{V}}^{+}\right).$ (5)
3.3 Negative Adaptation
Having introduced the positive branch of our DualAdapter, we now explore the negative inference capabilities of VLMs, adapting them to downstream tasks from a negative perspective. Specifically, we craft negative prompts and generate negative image prototypes to direct CLIP towards more confidently excluding incorrect classes based on the given input image.
Negative Textual Adapter. Recall that the positive textual cache, denoted as $\mathsf{T}_{\mathtt{cache}}^{+}$, is constructed based on the positive class descriptor prompt such as “A photo of a {CLASS}.” Conversely, we introduce a set of negative prompts (e.g., “A photo of no {CLASS},” “A photo without {CLASS}”) that convey meanings opposite in semantics [^1] The comprehensive collection of positive and negative prompts utilized for each dataset is available in the Supplementary Materials.. By leveraging the textual features $f_{t}^{-}$ derived from the negative text descriptions linked with these prompts, we again precompute a weight matrix and construct negative textual cache $\mathsf{T}_{\mathtt{cache}}^{-}=[f_{t_{1}}^{-}\,f_{t_{2}}^{-}\,\cdots\,f_{t_{C}}^{-}]^{\top}\in\mathbb{R}^{C\times d}$. Intuitively, if the feature of an input image $f_{\mathcal{I}}$ closely resembles the negative textual feature for a specific class $c$, it strongly suggests that the image does not belong to class $c$, i.e., $\mathbb{P}(y=c|\mathcal{I})\propto 1-f_{\mathcal{I}}^{\top}f_{t_{c}}^{-}$. Considering the zero-shot CLIP model is not designed to perform this kind of negative prediction, we incorporate a learnable residual $\mathcal{R}_{\mathsf{T}}^{-}\in\mathbb{R}^{C\times d}$ to refine the negative text embeddings through supervised task-specific training. The prediction logit is subsequently given by:
$\mathsf{T}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{T}}^{-}\right),\quad\mathcal{S}_{\mathsf{T}}^{-}=\delta_{\mathsf{T}}\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times C},$ (6)
where $\delta_{\mathsf{T}}$ is a fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{T}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{T}}^{+}$, and $\mathbf{1}\in\mathbb{R}^{1\times C}$ denotes an all-ones $C$-dimensional vector.
Negative Visual Adapter. In the visual domain, we also similarly pursue some negative image features to enable negative predictions. To achieve this, we naturally treat all $(C-1)K$ images from the other $C-1$ classes as negative samples for a given class $c$. These negative image samples hold the same property as the negative text descriptions: they are explicitly selected to be non-representative of a specific class, which means that the higher the similarity to them, the lower the probability that the image is classified to that class. To mitigate individual biases, we randomly select one image from each of the $C-1$ classes and compute the average of their extracted features to represent the negative image prototypes [56]. In this way, we can get a total of $K$ negative image prototypes for each of the $C$ classes, thereby constructing a negative visual cache $\mathsf{V}_{\mathtt{cache}}^{-}\in\mathbb{R}^{CK\times d}$. Similar to Eq. (4), we compute the affinities between the input image and those negative image prototypes, and calculate the classification logit:
$\mathcal{A}^{-}\!=\!\delta_{\mathsf{V}}\exp\left(-\beta\left(\mathbf{1}-\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\right)\right)$ $\!=\!\delta_{\mathsf{V}}\exp\left(-\beta f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times CK},$ (7)
$\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}L$ $\in\mathbb{R}^{1\times C}.$ (8)
where $\delta_{\mathsf{V}}$ is another fixed scaling parameter that adjusts $\mathcal{A}^{-}$/$\mathcal{S}_{\mathsf{V}}^{-}$ to match the mean value of $\mathcal{A}^{+}$/$\mathcal{S}_{\mathsf{V}}^{+}$, $L\in\mathbb{R}^{CK\times C}$ denotes the corresponding one-hot labels.
To further encode the diverse fine-grained visual knowledge associated with the specific downstream task, we introduce an extra set of learnable residual parameters, $\mathcal{R}_{\mathsf{V}}^{-}\in\mathbb{R}^{C\times d}$, which is similarly broadcast to refine the negative visual cache:
@@ 3.3 Negative Adaptation
$\mathsf{V}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{V}}^{-}\right).$ (9)
3.4 Unified Dual-Path Adapter
To effectively adapt VLMs to downstream tasks, our DualAdapter combines the functionalities of all four introduced adapters. As illustrated in Figure 3 (Left), we provide an overview of the specific example features stored in the cache for the class dog. By measuring the similarities between the input image feature and these cached example features, our DualAdapter produces a unified final prediction of the probability that the input image represents a dog. The same process is consistently applied across all classes to perform recognition tasks.
DualAdapter Inference. To derive the final classification scores, we aggregate the four outputs from dual-path (i.e., positive and negative) adapters spanning both textual and visual modalities, expressed as:
@@ 3.4 Unified Dual-Path Adapter
$\mathcal{S}_{\mathsf{final}}=\lambda\left(\mathcal{S}_{\mathsf{T}}^{+}+\mathcal{S}_{\mathsf{V}}^{+}\right)+(1-\lambda)\left(\mathcal{S}_{\mathsf{T}}^{-}+\mathcal{S}_{\mathsf{V}}^{-}\right).$ (10)
@@
Here, $\lambda$ serves as a tuning hyper-parameter to balance the contribution of positive and negative adapter logits. Throughout the training process, the collection of learnable parameters =$\mathcal{R}=\{\mathcal{R}_{\mathsf{T}}^{+},\mathcal{R}_{\mathsf{V}}^{+},\mathcal{R}_{\mathsf{T}}^{-},\mathcal{R}_{\mathsf{V}}^{-}\}$ is updated via stochastic gradient descent, employing a cross-entropy loss function for optimization.
Figure 3: Visualization of cosine similarities on the validation set of ImageNet [4]. (Left) Example features of the class dog stored in the cache; (Middle) Distribution of pairwise image-text similarities with dual-path features in textual adapters; (Right) Distribution of pairwise image-image similarities with dual-path features in visual adapters.
@@
Effectiveness of Four Adapters. Fig. 3 qualitatively shows the effectiveness of the four adapters we designed. In the middle and right sub-figure, we visualize the distributions of the pairwise cosine similarities between the features of the input image and the textual/visual features stored in the adapters on ImageNet [4] validation set. From Fig. 3, we can observe the following statistical patterns: (1) As we expected, the input image feature is more similar with positive features from the same class (blue $>$ orange) and negative features from other classes (red $>$ green) across both two modalities; (2) Within the textual modality, negative features tend to be less similar to the input image compared to positive features (red $<$ blue, green $<$ orange), since the negative prompts are less common in CLIP [48] training corpus; (3) Within the visual modality, the similarity distribution of negative features occupies an intermediate position between the positive features of the same and different classes (orange $<$ green/red $<$ blue). This is because the negative visual features are constructed by averaging the image features across all other classes, inherently leading to a more generic representation that lacks the distinctiveness characteristic of positive features within a single class.
3.5 Similarity-Based Label Refinement
Previous adapter-style fine-tuning methods [79, 64] generally assume that the few-shot training samples are meticulously curated to provide accurate representation for each class. However, in practical scenarios, this assumption often does not hold true. We recognize that not every image is of high quality or equally representative of its respective class, which can be particularly detrimental in the context of few-shot learning, where the availability of training images is extremely limited. For example, we provide the t-SNE [38] visualization of few-shot image features from the OxfordPets [45] dataset in Fig. 4 (Left). We argue that the few-shot samples are highly noisy: there is even a data point from the Basset Hound class that overlaps with the prototype (i.e., average feature) of Cocker Spaniel.
In response to this, we devise an unsupervised similarity-based label refinement process to address this challenge. Recall that in our visual adapter, we calculate the image-image affinities between the input image and few-shot training images, with all few-shot images being assigned equal weights [^2] Each few-shot image is assigned a one-hot label, corresponding to a weight of 1. as detailed in Eq. (4). As we discussed, this approach is possibly problematic in the noisy few-shot learning setting. Consequently, we extend the visual adapter to assign non-uniform confidence scores according to pairwise similarities, effectively downweighting outliers and emphasizing more representative samples.
@@ 3.5 Similarity-Based Label Refinement
Specifically, our refinement process is based on an intuitive assumption: the representative image feature is closer to other image features from the same class than those low-quality outliers [26]. Based on this assumption, given the $K$-shot image features $\{f_{v_{c}}^{(i)}\}_{i=1}^{K}$ from a specific class $c$, we calculate the average cosine similarities of each image feature to others:
@@
We then assign non-uniform confidence scores $w_{c}^{(i)}$ and compute refined label values $\ell_{c}^{(i)}$ for all $K$-shot image features based on their similarities to others:
@@
where $\tau$ is a temperature hyper-parameter to control the intensity of our refinement. In particular, as $\tau\rightarrow\infty$, our refinement becomes inactive, assigning a label value of 1 to all instances. Conversely, as $\tau\rightarrow 0$, our refinement process becomes more selective, assigning the highest label value exclusively to the most representative image feature, while all other features are assigned 0.
This refinement process is applied across each class, wherein the original one-hot labels $L$, are weighted by the new label values computed in Eq. (12), resulting in the updated labels $\mathbb{L}$. This refinement process is incorporated into both the positive and negative visual adapters; therefore we can rewrite Eqs. (4) and (8) as:
$\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}\mathbb{L}\in\mathbb{R}^{1\times C},\quad\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}\mathbb{L}\in\mathbb{R}^{1\times C}.$ (13)
Figure 4: Motivation for the proposed similarity-based label refinement. (Left) t-SNE [38] visualization of visual features for 4 representative classes from the OxfordPets [45] dataset, with prototype [56] of each class denoted by the crossmark $\times$; (Middle) Our refined labels for 4 example images from class Beagle; (Right) An illustrative plot of the modeled distributions using one-hot labels and our refined labels for the Beagle class.
Effectiveness of Label Refinement. In Fig. 4 (Middle), we present our refined label values for 4 example images from class Beagle in OxfordPets [45] dataset. Notably, our unsupervised refinement approach functions by measuring the pairwise similarities among 8-shot images from class Beagle, without requiring any extra parameters or additional training. Our similarity-based refinement process allocates higher values to the more representative samples at the top, and lower values to the lower-quality images at the bottom, where the Beagle is either obscured or presented in black and white. In Fig. 4 (Right), we provide an illustration of the modeled visual feature distribution for the Beagle class using one-hot labels and our refined labels. Through our refinement process, we are better equipped to capture the underlying feature distribution (typically, a normal distribution [68, 81]) of the Beagle class, whereas the use of one-hot labels restricts us to modeling merely a uniform distribution.

### v2

@@ 3.3 Learning the Positive Classifier
Positive Textual Branch. As shown in Eq. (4), CLIP can make zero-shot predictions utilizing a textual cache, which stores the textual features from positive text descriptions (e.g., “A photo of a  {CLASS}”). To avoid ambiguity, we denote this cache as $\mathsf{T}_{\mathtt{cache}}^{+}=[f_{t_{1}}^{+}\,f_{t_{2}}^{+}\,\cdots\,f_{t_{C}}^{+}]^{\top}\in\mathbb{R}^{C\times d}$. With a small set of annotated training images, we can update the textual features by introducing a group of learnable residual parameters $\mathcal{R}_{\mathsf{T}}^{+}\in\mathbb{R}^{C\times d}$ to integrate task-specific knowledge:
@@
$\mathsf{T}_{\mathtt{cache}}^{+}\leftarrow\mathtt{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{T}}^{+}\right),$
$\mathcal{S}_{\mathsf{T}}^{+}=f_{v}\mathsf{T}_{\mathtt{cache}}^{+\quad\top}\in\mathbb{R}^{C}.$ (5)
Here, $\mathtt{Normalize}$ denotes the L2-normalization applied to each row of the matrix.
Positive Visual Branch. We further extend the recognition capability of the CLIP model by constructing a visual cache-based classifier, which operates by measuring image-image similarities between the input image feature and few-shot image features. Specifically, we store all $CK$ image features in a precomputed visual cache, denoted as $\mathsf{V}_{\mathtt{cache}}^{+}=[f_{v_{1}}^{(1)}\,f_{v_{1}}^{(2)}\,\cdots\,f_{v_{C}}^{(K)}]^{\top}\in\mathbb{R}^{CK\times d}$. Their one-hot labels are also correspondingly recorded in $L\in\mathbb{R}^{CK\times C}$. Given an image feature $f_{v}$ to be classified, we calculate its image-image affinities and obtain the logits $\mathcal{S}_{\mathsf{V}}^{+}$:
@@
$\mathsf{V}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{V}}^{+}\right),$
@@
$\mathcal{S}_{\mathsf{V}}^{+}=\mathcal{A}\left(f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)L\in\mathbb{R}^{C},$ (6)
where $\mathcal{A}\left(f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)=\alpha\exp\left(-\beta\left(1-f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\right)$ calculates the affinity, $\alpha$ represents a balance scalar and $\beta$ denotes a modulating parameter. Learnable residue $\mathcal{R}_{\mathsf{V}}^{+}\in\mathbb{R}^{C\times d}$ is also introduced, which is broadcast to $\mathbb{R}^{CK\times d}$ and added to the visual cache to refine the visual features.
3.4 Learning the Negative Classifier
Having introduced the positive classifier, we now explore constructing and learning a negative dissimilarity-based classifier, which directs CLIP towards more confidently excluding incorrect classes based on the given input image. Conceptually, our goal is to mine a general negative feature for each class $c$, which is absent in samples from class $c$ but present in samples from all other classes.
Negative Textual Branch. Recall that the positive textual cache, denoted as $\mathsf{T}_{\mathtt{cache}}^{+}$, is constructed based on the positive class descriptor prompt such as “A photo of a {CLASS}.” Conversely, we introduce a set of negative prompts (e.g., “A photo of not a {CLASS},” “A photo without {CLASS}”) that convey opposite semantics meanings. By leveraging the textual features $f_{t}^{-}$ derived from the negative text descriptions linked with these prompts, we again precompute a weight matrix and construct negative textual cache $\mathsf{T}_{\mathtt{cache}}^{-}=[f_{t_{1}}^{-}\,f_{t_{2}}^{-}\,\cdots\,f_{t_{C}}^{-}]^{\top}\in\mathbb{R}^{C\times d}$. Intuitively, if the feature of an input image $f_{v}$ closely resembles the negative textual feature for a specific class $c$, it strongly suggests that the image does not belong to class $c$, i.e., $\mathbb{P}(y=c|\boldsymbol{x})\propto 1-f_{v}^{\top}f_{t_{c}}^{-}$. Through supervised task-specific training, we optimize a learnable residue $\mathcal{R}_{\mathsf{T}}^{-}\in\mathbb{R}^{C\times d}$ and obtain the negative predictions based on dissimilarities $1-f_{v}\mathsf{T}_{\mathtt{cache}}^{-}$:
$\mathsf{T}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{T}}^{-}\right),$
$\mathcal{S}_{\mathsf{T}}^{-}=\delta_{\mathsf{T}}\left(1-f_{v}\mathsf{T}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{C},$ (7)
where $\delta_{\mathsf{T}}$ is a fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{T}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{T}}^{+}$.
Figure 3: Visualization of cosine similarities on ImageNet [6] validation set. We present distributions of pairwise similarities between the input image feature and both the learned positive and negative features from textual (Left) and visual (Right) modalities.
Negative Visual Branch. In the visual domain, we also similarly pursue some negative image features to be non-representative of a specific class, which means that the higher the similarity to them, the lower the probability that the image is classified to that class, i.e., $\mathbb{P}(y=c|\boldsymbol{x})\propto 1-f_{v}^{\top}f_{v_{c}}^{-}$. To achieve this, we randomly select one image from each of the $C-1$ classes and compute the average of their extracted features to represent the negative features. In this way, we can get a total of $K$ negative visual features for each of the $C$ classes, thereby constructing a negative visual cache $\mathsf{V}_{\mathtt{cache}}^{-}\in\mathbb{R}^{CK\times d}$. Symmetric to Eq. (6), we compute the negative affinities to obtain the classification logits:
$\mathsf{V}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{V}}^{-}\right),$
$\mathcal{S}_{\mathsf{V}}^{-}=\delta_{\mathsf{V}}\mathcal{A}\left(1-f_{v}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)L\in\mathbb{R}^{C}.$ (8)
where $\delta_{\mathsf{V}}$ is another fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{V}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{V}}^{+}$. Here, we also introduce a set of learnable parameters $\mathcal{R}_{\mathsf{V}}^{-}\in\mathbb{R}^{C\times d}$ to refine the negative visual features.
Final Inference. As discussed in Section 3.1, we ensemble the predictions from both classifiers to derive the final classification scores and predictions:
$\mathcal{S}_{\mathsf{final}}=\lambda\left(\mathcal{S}_{\mathsf{T}}^{+}+\mathcal{S}_{\mathsf{V}}^{+}\right)+(1-\lambda)\left(\mathcal{S}_{\mathsf{T}}^{-}+\mathcal{S}_{\mathsf{V}}^{-}\right),$
$\mathbb{P}(y|\boldsymbol{x})=\mathtt{Softmax}\left(\mathcal{S}_{\mathsf{final}}\right).$ (9)
@@ 3.4 Learning the Negative Classifier
Here, $\lambda$ serves as a tuning hyper-parameter to balance the contribution of positive and negative logits. Throughout the training process, the collection of learnable parameters =$\mathcal{R}=\{\mathcal{R}_{\mathsf{T}}^{+},\mathcal{R}_{\mathsf{V}}^{+},\mathcal{R}_{\mathsf{T}}^{-},\mathcal{R}_{\mathsf{V}}^{-}\}$ is updated via stochastic gradient descent with a cross-entropy loss.
@@
Qualitative Visualizations. Figure 3 qualitatively shows the effectiveness of both the positive and negative classifiers we designed. Specifically, we visualize the distributions of pairwise cosine similarities between the input image feature $f_{v}$ and both the learned positive and negative features from two modalities ($f_{t}^{+},f_{t}^{-},f_{v}^{+},f_{v}^{-}$) on the validation set of ImageNet [6]. We can observe the following statistical patterns: (1) As we expected, the input image feature is more similar with positive features from the same class (blue $>$ orange) and negative features from other classes (red $>$ green) across both two modalities; (2) Within the visual modality, the similarity distribution of negative features occupies an intermediate position between the positive features of the same and different classes (orange $<$ green/red $<$ blue). This is because the negative visual features are constructed by averaging the image features across all other classes, inherently leading to a more generic representation that lacks the distinctiveness characteristic of positive features within a single class; (3) Within the textual modality, negative features tend to be less similar to the input image compared to positive features (red $<$ blue, green $<$ orange), since the negative prompts are less common in CLIP [38] training corpus.
3.5 Few-Shot Instance Reweighting
In the few-shot adaptation setting, our classifier is trained using only a limited number of samples, with each sample making a significant contribution to the formation of the final decision boundary. Consequently, our VLMs are particularly vulnerable to potential noise in the few-shot sample set. In Figure 4 (Left), we simulate real-world noise scenarios by randomly flipping the labels of a portion of support samples. We demonstrate that the performance of Tip-Adapter-F [71] drastically decreases from 65.52% to 62.47% when the labels of 8 out of the 16 samples per class are randomly flipped. Moreover, even if the labels are all correct, we recognize that not every image is of high quality or equally representative of its respective class, as shown in Figure 4 (Right). To address this issue, we have developed a few-shot instance reweighting technique to assign non-uniform confidence scores to each sample, effectively downweighting outliers (or mislabelled samples) and prioritizing more representative samples.
Figure 4: Few-shot instance reweighting. (Left) The performance of Tip-Adapter-F [71] degrades drastically when label noise exists in the few-shot sample set; (Right) t-SNE [53] visualization of visual features for 4 random classes from the OxfordPets [35] dataset, where some outliers are marked with red circles.
@@ 3.5 Few-Shot Instance Reweighting
Specifically, our proposed instance reweighting is based on an intuitive assumption: the representative image feature is closer to other image features from the same class than those low-quality outliers. Based on this assumption, given the $K$-shot image features $\{f_{v_{c}}^{(i)}\}_{i=1}^{K}$ from a specific class $c$, we calculate the average cosine similarities of each image feature to others:
@@
We then assign non-uniform weights $w_{c}^{(i)}$ and compute the reweighted confidences $\ell_{c}^{(i)}$ for all $K$-shot image features based on their average similarities to others:
@@
where $\tau$ is a temperature hyper-parameter to control the intensity of our instance reweighting.
This process is applied across each class, wherein the original one-hot labels $L$, are reweighted using the new confidence values calculated in Eq. (11) to yield $\mathbb{L}$. By incorporating this reweighting technique into the visual branches of both classifiers, we can reformulate Eqs. (6) and (8) as follows:
$\mathcal{S}_{\mathsf{V}}^{+}=\mathcal{A}\left(f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\mathbb{L}^{+},\,\mathcal{S}_{\mathsf{V}}^{-}=\delta_{\mathsf{V}}\mathcal{A}\left(1-f_{v}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\mathbb{L}^{-}.$ (12)